# Large Language Models for Scientific Idea Generation: A Creativity-Centered Survey

### v1

@@ 7.3 LLM-as-a-Judge Evaluation
The LLM-as-a-Judge paradigm reframes large language models from generators to evaluators capable of approximating human reasoning. Early benchmarks such as MT-Bench and Chatbot Arena  (zheng2023judgingllmasajudgemtbenchchatbot) showed that LLMs can reliably proxy human preferences and judgments even in zero-shot settings, establishing their use as scalable evaluators across domains. In scientific discovery, LLM-judges are now applied to assess plausibility, novelty, and impact of hypotheses, occupying a practical middle ground between purely computational metrics (which often miss conceptual depth), execution-based validation (which can be costly or infeasible), and labor-intensive human review.
@@ 7.3.1 Prompt-based / Inference-time judges
LLMs’ zero- and few-shot in-context reasoning enables them to act as first-line evaluators with minimal supervision. In the zero-shot setting, models are guided by reviewing rubrics encoded directly in prompts. For example, the GPT-4o-based AI Scientist (lu2024aiscientistfullyautomated; yamada2025aiscientistv2) applies NeurIPS-style reviewing criteria to generate structured assessments of generated manuscripts. LiveIdeaBench (ruan2025liveideabenchevaluatingllmsdivergent) employs a multi-agent jury of LLM evaluators—each assigned distinct system prompts and roles—to rate generated ideas across originality, feasibility, fluency, flexibility, and significance. ResearchAgent (baek2024researchagent) coordinates multiple reviewer agents to iteratively critique and refine hypotheses. Similarly, Chain-of-Idea (li2024chain) introduces a comparative “idea arena,” evaluating pairs of hypotheses against each other rather than assigning absolute scores to reduce judgment drift and improve calibration.
Beyond pure prompting, a large body of work augments these evaluators with retrieval pipelines that ground LLM judgments in empirical evidence. Such retrieval-based inference enhances reliability in metrics like novelty, feasibility, and impact by contextualizing model reasoning with recent literature. For instance, the Idea Novelty Checker (shahid2025literaturegroundednoveltyassessmentscientific) retrieves semantically related papers via dense embeddings, filters and re-ranks them with LLM reasoning across facets such as purpose and mechanism, achieving over 13% higher agreement with expert judgments. DeepReview (zhu2025deepreview) extends this approach with multi-step retrieval and synthesis stages, generating structured “novelty verification” reports that summarize evidence gaps and methodological contributions.
Retrieval grounding has similarly advanced feasibility assessment. Matter-of-Fact (jansen2025matteroffactbenchmarkverifyingfeasibility) benchmarks literature-grounded reasoning in materials science, testing whether claims are feasible given pre-publication data. SciPIP (wang2024scipip) integrates citation-graph reasoning and semantic retrieval to estimate methodological plausibility and resource constraints, while IdeaSynth (pu2025ideasynth) decomposes ideas into problem–method–evaluation triplets and measures proximity to historically successful research, estimating both feasibility and experimental cost.
Finally, works like aiXiv (zhang2025aixivnextgenerationopenaccess) demonstrate large-scale retrieval-augmented evaluation pipelines for impact prediction, combining literature grounding with structured LLM judgment.
@@ 7.3.2 Supervised fine-tuned judges
A more robust direction uses SFT on large review datasets, aligning LLMs to human scoring behavior. Here, models are exposed to richly annotated review datasets containing both scores and rationales, enabling them to internalize reviewing patterns. CycleReviewer (weng2025cycleresearcher), trained as a generative reward model on the Review-5K dataset, achieves a decision accuracy of 74.24%, complementing human expertise with consistent automated scoring. Building on this, OpenReviewer idahl-ahmadi-2025-openreviewer applies full-parameter SFT to a Llama-3.1-8B-Instruct model using the 80K reviews from top conferences, producing structured reviews with strong reasoning quality.
DeepReview (zhu2025deepreview) pushes SFT-based evaluation further through a significantly larger and more structured training corpus—DeepReview-13K—which integrates full research papers, multi-stage reasoning traces, and final human-aligned assessments. During inference, it employs retrieval-augmented review synthesis, self-verification, and reflection modules that enable iterative refinement of review reasoning. The resulting 14B-parameter model, despite its moderate size, surpasses larger models like CycleReviewer-70B, GPT-o1 and Deepseek-R1 by dynamically scaling inference complexity, demonstrating that dataset design and structured supervision can outweigh pure model scale.
These SFT pipelines emphasize the importance of dataset curation and training objectives: the models learn scoring intuitions from reviewer’ annotations, which yields high correlation with human judgments. However, recent studies highlight several limitations of SFT in capturing scientific novelty. (shuieh2025assessingrobustnessspuriouscorrelations) show that SFT often induces overreliance on spurious cues in the data, reducing robustness when those features fail.
@@ 7.3.3 Reinforcement-optimized judges
The third regime advances beyond SFT memorization toward generalization via RL (chu2025sftmemorizesrlgeneralizes). While SFT-based reviewers mimic human scoring behavior, they often overfit to annotation bias or linguistic patterns, limiting their ability to assess novel or out-of-distribution ideas. RL optimization instead trains judges to reason through evaluation criteria, promoting deeper alignment, calibration, and consistency.
Recent frameworks such as J1 (whitehouse2025j1) show that tailored RL objectives can internalize reasoning depth rather than surface agreement. J1 fine-tunes reviewers using reward signals from rubric adherence and multi-step self-consistency, outperforming SFT and DPO baselines in cross-domain generalization and inter-rater reliability. Similarly, ReviewRL (zeng2025reviewrl) warm-starts from SFT and applies Reinforce++ optimization for review quality and stability, achieving superior coherence, factuality, and fairness across venues like NeurIPS and ICLR.
REMOR (taechoyotin2025remorautomatedpeerreview) extends this direction with multi-objective RL that balances factual accuracy, reasoning depth, and creativity encouragement. By coupling review text with reasoning traces and composite rewards—covering alignment, novelty, and consistency—it improves citation prediction accuracy and reduces verbosity compared to SFT-only baselines.
Finally, HARPA (vasu2025harpatestabilitydrivenliteraturegroundedframework) adapts the reinforcement-optimized paradigm to hypothesis generation. Its learned reward model scores ideas by testability and groundedness using prior experimental outcomes as feedback. HARPA-generated hypotheses outperform strong AI-researcher baselines in feasibility (+0.78) and groundedness (+0.85) on human ratings, achieving higher real-world execution success in the CodeScientist pipeline (20 vs. 11 successful runs). This illustrates how RL-trained evaluators can not only assess but also guide ideation toward scientifically meaningful discoveries. By directly optimizing for rating stability and agreement with human gold standards, RL-trained critics achieve more reliable judgments across domains.
@@ 7.3.4 Hybrid Frameworks
To overcome these constraints, hybrid evaluators integrate learned judgment policies with test-time reasoning and external tool use. DeepReviewer (zhu2025deepreview) exemplifies the fusion of training-based alignment and inference-time adaptability. It combines the structured evaluative intuition learned during supervised fine-tuning with external retrieval and self-verification mechanisms activated at inference, allowing the model to ground its reasoning dynamically in relevant literature. Through its scalable inference modes (Fast, Standard, Best), DeepReviewer-14B adjusts deliberation depth to match computational budgets, maintaining both efficiency and analytical rigor. This synergy between training priors and adaptive test-time reasoning enables DeepReviewer to outperform much larger models such as CycleReviewer-70B, producing richer and more accurate reviews with fewer tokens and strong resilience to adversarial perturbations—despite the absence of explicit robustness training. ReviewRL (zeng2025reviewrl) similarly augments its RL-trained critic with retrieval-augmented context (e.g., ArXivMCP) to ground evaluations in verifiable scientific evidence. Such hybrid systems balance the efficiency of automation with the rigor of human-like review, highlighting a consensus that future evaluation frameworks will need to fuse training-based alignment with external grounding to achieve both reliability and transparency.
@@
Recent work comparing human and LLM-based reviewers highlights complementary strengths but clear gaps. Zero-shot systems such as Claude 3.5 Sonnet produce structured, comprehensive reviews yet lack the contextual depth and domain reasoning of humans (Renata2025AIRA). Similarly, GPT-4o performs well in summarizing contributions but misses weaknesses and critical insights (Li2025UnveilingTM). Broader evaluations show that while GPT-based models reach over 60% accuracy in review prediction, they remain limited in long-context and analytical reasoning (Zhou2024IsLA). Other analyses further warn that LLMs, though reducing reviewer fatigue, risk amplifying bias and diminishing accountability(Hosseini2023FightingRF). Overall, LLM-based judges provide scalable and consistent first-pass evaluations but still require human calibration to ensure fairness, nuance, and contextual rigor (Table 4).
@@ 7.4 Challenges in Evaluation
Among the four Ps of creativity, the Product dimension is uniquely challenging: it represents the measurable outcome of creativity, whereas Person, Process, and Press serve as sources that shape how such products emerge (kozbelt2010theories; 4pCreativityModel). For scientific idea generation, this means that evaluation is not just an auxiliary step but the central bottleneck—where questions of novelty, feasibility, diversity, and impact must ultimately be resolved. Yet current practices reveal deep limitations that hinder progress.
First, evaluation suffers from subjectivity and inconsistency. Human raters often diverge in their judgments depending on expertise, background, or expectations, and are prone to systematic biases—for instance, underrating AI-generated ideas or devaluing content once labeled as machine-produced (glikson2020humantrust; Langham2022GenderBias; zhu2025humanbiasfaceai). Automated proxies such as embedding similarity or LLM critics are not immune either, as they may replicate biases encoded in training data or drift with prompting (laskar2024systematicsurveycriticalreview; wang2025enablingaiscientistsrecognize). A second challenge is the lack of standardization: most studies define their own evaluation metrics, leading to fragmented practices and results that are not reproducible or comparable across benchmarks (gehrmann2022repairingcrackedfoundationsurvey; laskar2024systematicsurveycriticalreview). Relatedly, many pipelines rely on short-term proxies. Metrics such as semantic dispersion or overlap with retrieved papers capture immediate surface features, but true scientific impact unfolds over years and depends on contextual uptake (ismayilzada2025creativityaiprogresseschallenges). Fourth, there is the risk of bias amplification. Both human reviewers and LLM-based judges can propagate historical inequities, undervaluing unconventional or marginalized research directions unless carefully calibrated (gehrmann2022repairingcrackedfoundationsurvey). Finally, a persistent scalability gap remains unresolved: while expert review provides nuance and reliability, it is expensive and prone to fatigue effects, whereas automated metrics scale efficiently but fail to capture the depth of human judgment (thelwall2025researchqualityevaluationai).

### v2

@@ 7.3 LLM-as-a-Judge Evaluation
The LLM-as-a-Judge paradigm reframes large language models from generators to evaluators capable of approximating aspects of human evaluative judgment. Early benchmarks such as MT-Bench and Chatbot Arena  (Zheng et al., 2023) showed that LLMs can reliably proxy human preferences and judgments even in zero-shot settings, establishing their use as scalable evaluators across domains. In scientific discovery, LLM-judges are now applied to assess plausibility, novelty, and impact of hypotheses, occupying a practical middle ground between purely computational metrics (which often miss conceptual depth), execution-based validation (which can be costly or infeasible), and labor-intensive human review. However, these systems do not replicate domain expertise or epistemic reasoning and should be interpreted as preference-aligned evaluators rather than substitutes for expert judgment.
@@ 7.3.1 Prompt-based / Inference-time judges
LLMs’ zero- and few-shot in-context reasoning enables them to act as first-line evaluators with minimal supervision. In the zero-shot setting, models are guided by reviewing rubrics encoded directly in prompts. For example, the GPT-4o-based AI Scientist (Lu et al., 2024a; Yamada et al., 2025) applies NeurIPS-style reviewing criteria to generate structured assessments of generated manuscripts. LiveIdeaBench (Ruan et al., 2024) employs a multi-agent jury of LLM evaluators—each assigned distinct system prompts and roles—to rate generated ideas across originality, feasibility, fluency, flexibility, and significance. ResearchAgent (Baek et al., 2024) coordinates multiple reviewer agents to iteratively critique and refine hypotheses. Similarly, Chain-of-Idea (Li et al., 2024a) introduces a comparative “idea arena,” evaluating pairs of hypotheses against each other rather than assigning absolute scores to reduce judgment drift and improve calibration.
Beyond pure prompting, a large body of work augments these evaluators with retrieval pipelines that ground LLM judgments in empirical evidence. Such retrieval-based inference enhances reliability in metrics like novelty, feasibility, and impact by contextualizing model reasoning with recent literature. For instance, the Idea Novelty Checker (Shahid et al., 2025) retrieves semantically related papers via dense embeddings, filters and re-ranks them with LLM reasoning across facets such as purpose and mechanism, achieving over 13% higher agreement with expert judgments. DeepReview (Zhu et al., 2025b) extends this approach with multi-step retrieval and synthesis stages, generating structured “novelty verification” reports that summarize evidence gaps and methodological contributions.
Retrieval grounding has similarly advanced feasibility assessment. Matter-of-Fact (Jansen et al., 2025) benchmarks literature-grounded reasoning in materials science, testing whether claims are feasible given pre-publication data. SciPIP (Wang et al., 2024c) integrates citation-graph reasoning and semantic retrieval to estimate methodological plausibility and resource constraints, while IdeaSynth (Pu et al., 2025) decomposes ideas into problem–method–evaluation triplets and measures proximity to historically successful research, estimating both feasibility and experimental cost.
Finally, works like aiXiv (Zhang et al., 2025b) demonstrate large-scale retrieval-augmented evaluation pipelines for impact prediction, combining literature grounding with structured LLM judgment.
@@ 7.3.2 Supervised fine-tuned judges
A more robust direction uses SFT on large review datasets, aligning LLMs to human scoring behavior. Here, models are exposed to richly annotated review datasets containing both scores and rationales, enabling them to internalize reviewing patterns. CycleReviewer (Weng et al., 2025a), trained as a generative reward model on the Review-5K dataset, achieves a decision accuracy of 74.24%, complementing human expertise with consistent automated scoring. Building on this, OpenReviewer Idahl and Ahmadi (2025) applies full-parameter SFT to a Llama-3.1-8B-Instruct model using the 80K reviews from top conferences, producing structured reviews with strong reasoning quality.
DeepReview (Zhu et al., 2025b) pushes SFT-based evaluation further through a significantly larger and more structured training corpus—DeepReview-13K—which integrates full research papers, multi-stage reasoning traces, and final human-aligned assessments. During inference, it employs retrieval-augmented review synthesis, self-verification, and reflection modules that enable iterative refinement of review reasoning. The resulting 14B-parameter model, despite its moderate size, surpasses larger models like CycleReviewer-70B, GPT-o1 and Deepseek-R1 by dynamically scaling inference complexity, demonstrating that dataset design and structured supervision can outweigh pure model scale.
These SFT pipelines emphasize the importance of dataset curation and training objectives: the models learn scoring intuitions from reviewer’ annotations, which yields high correlation with human judgments. However, recent studies highlight several limitations of SFT in capturing scientific novelty. (Shuieh et al., 2025) show that SFT often induces overreliance on spurious cues in the data, reducing robustness when those features fail.
@@ 7.3.3 Reinforcement-optimized judges
The third regime advances beyond SFT memorization toward generalization via RL (Chu et al., 2025). While SFT-based reviewers mimic human scoring behavior, they often overfit to annotation bias or linguistic patterns, limiting their ability to assess novel or out-of-distribution ideas. RL optimization instead trains judges to reason through evaluation criteria, promoting deeper alignment, calibration, and consistency.
Recent frameworks such as J1 (Whitehouse et al., 2025) show that tailored RL objectives can internalize reasoning depth rather than surface agreement. J1 fine-tunes reviewers using reward signals from rubric adherence and multi-step self-consistency, outperforming SFT and DPO baselines in cross-domain generalization and inter-rater reliability. Similarly, ReviewRL (Zeng et al., 2025) warm-starts from SFT and applies Reinforce++ optimization for review quality and stability, achieving superior coherence, factuality, and fairness across venues like NeurIPS and ICLR.
REMOR (Taechoyotin and Acuna, 2025) extends this direction with multi-objective RL that balances factual accuracy, reasoning depth, and creativity encouragement. By coupling review text with reasoning traces and composite rewards—covering alignment, novelty, and consistency—it improves citation prediction accuracy and reduces verbosity compared to SFT-only baselines.
Finally, HARPA (Vasu et al., 2025) adapts the reinforcement-optimized paradigm to hypothesis generation. Its learned reward model scores ideas by testability and groundedness using prior experimental outcomes as feedback. HARPA-generated hypotheses outperform strong AI-researcher baselines in feasibility (+0.78) and groundedness (+0.85) on human ratings, achieving higher real-world execution success in the CodeScientist pipeline (20 vs. 11 successful runs). This illustrates how RL-trained evaluators can not only assess but also guide ideation toward scientifically meaningful discoveries. By directly optimizing for rating stability and agreement with human gold standards, RL-trained critics achieve more reliable judgments across domains.
@@ 7.3.4 Hybrid Frameworks
To overcome these constraints, hybrid evaluators integrate learned judgment policies with test-time reasoning and external tool use. DeepReviewer (Zhu et al., 2025b) exemplifies the fusion of training-based alignment and inference-time adaptability. It combines the structured evaluative intuition learned during supervised fine-tuning with external retrieval and self-verification mechanisms activated at inference, allowing the model to ground its reasoning dynamically in relevant literature. Through its scalable inference modes (Fast, Standard, Best), DeepReviewer-14B adjusts deliberation depth to match computational budgets, maintaining both efficiency and analytical rigor. This synergy between training priors and adaptive test-time reasoning enables DeepReviewer to outperform much larger models such as CycleReviewer-70B, producing richer and more accurate reviews with fewer tokens and strong resilience to adversarial perturbations—despite the absence of explicit robustness training. ReviewRL (Zeng et al., 2025) similarly augments its RL-trained critic with retrieval-augmented context (e.g., ArXivMCP) to ground evaluations in verifiable scientific evidence. Such hybrid systems balance the efficiency of automation with the rigor of human-like review, highlighting a consensus that future evaluation frameworks will need to fuse training-based alignment with external grounding to achieve both reliability and transparency.
@@
Recent work comparing human and LLM-based reviewers highlights complementary strengths but clear gaps. Zero-shot systems such as Claude 3.5 Sonnet produce structured, comprehensive reviews yet lack the contextual depth and domain reasoning of humans (Renata and Lee, 2025). Similarly, GPT-4o performs well in summarizing contributions but misses weaknesses and critical insights (Li et al., 2025b). Broader evaluations show that while GPT-based models reach over 60% accuracy in review prediction, they remain limited in long-context and analytical reasoning (Zhou et al., 2024). Other analyses further warn that LLMs, though reducing reviewer fatigue, risk amplifying bias and diminishing accountability(Hosseini and Horbach, 2023). Overall, LLM-based judges provide scalable and consistent first-pass evaluations but still require human calibration to ensure fairness, nuance, and contextual rigor (Table 5).
Table 6: Scientific idea generation benchmarks and how they operationalize creativity dimensions.
Benchmark Task / Input Evaluation protocol Key strengths / limitations Creativity dimensions
Research-Bench Liu et al. • Inspiration retrieval • Idea / hypothesis composition • Ranking stage • Stage-wise evaluation • Ranking accuracy • Grounding • Human / LLM judges + End-to-end discovery modeling + Explicit grounding - Implicit scores Novelty, Grounding
IdeaBench  (Guo et al., 2024) • Research ideas from paper context • Title / abstract / references • Expert or LLM judges • Likert / pairwise ranking + Explicit ideation focus + Scalable templates - Qualitative feasibility - Judge sensitivity Novelty, Feasibility
LiveIdea-Bench  (Ruan et al., 2024) • Minimal context prompts • Keyword / short input • Divergent ideation test • Panel LLM judges • Multi-axis scoring • Score aggregation + Low-context creativity probe + Highly scalable - Judge bias / drift - Weak feasibility grounding Novelty, Diversity, Surprise
HypoBench  (Liu et al., 2025c) • Data-driven hypotheses • Real + synthetic tasks • Known structure • Ground-truth recovery • Discovery rate metrics • Human / LLM judges + Strong validity signal + Quantitative evaluation - Statistical focus - Limited ideation Validity, Testability
MatDesign  (Kumbhar et al., 2025) • Materials science domain • Knowledge-constrained prompts • Domain realism • Expert-aligned criteria • Feasibility, plausibility • Impact, testability + Rich feasibility checks + Realistic hypotheses - Narrow domain - High expert cost Novelty, Feasibility, Testability
Research-CodeBench  (Hua et al., 2025b) • Code-level realization • Implement recent methods • Execution-based tasks • Compile / run tests • Functional correctness • Quality checks + Objective feasibility signal + Execution grounded - No novelty measure - Realization $\neq$ ideation Feasibility, Correctness
LLM-SRBench  2025 • Equation discovery • Synthetic transformations • Contam.-aware design • Exact correctness • Fit accuracy • Robustness tests + Formal validity + Strong testability - Narrow discovery form - Limited ideation scope Validity, Testability
Matter-ofFact  Jansen et al. • Literature-supported claims • Materials science domain • Claim feasibility verification • Consistency checks • Human / LLM judges + Explicit feasibility grounding - Not generative by itself - Domain-specific Feasibility, Validity
@@ 7.4 Challenges in Evaluation
Among the four Ps of creativity, the Product dimension is uniquely challenging: it represents the measurable outcome of creativity, whereas Person, Process, and Press serve as sources that shape how such products emerge (Kozbelt et al., 2010; Gruszka and Tang, 2017). For scientific idea generation, this means that evaluation is not just an auxiliary step but the central bottleneck—where questions of novelty, feasibility, diversity, and impact must ultimately be resolved. Yet current practices reveal deep limitations that hinder progress.
First, evaluation suffers from subjectivity and inconsistency. Human raters often diverge in their judgments depending on expertise, background, or expectations, and are prone to systematic biases—for instance, underrating AI-generated ideas or devaluing content once labeled as machine-produced (Glikson and Woolley, 2020; Langham et al., 2022; Zhu et al., 2025c). Automated proxies such as embedding similarity or LLM critics are not immune either, as they may replicate biases encoded in training data or drift with prompting (Laskar et al., 2024; Wang et al., 2025e).
A second challenge is the lack of standardization: most studies define their own evaluation metrics, leading to fragmented practices and results that are not reproducible or comparable across benchmarks (Gehrmann et al., 2022; Laskar et al., 2024). Table 6 illustrates this heterogeneity: benchmarks vary not only in task framing (idea vs. hypothesis generation vs. discovery vs. code realization), but also in evaluation protocols (expert panels, LLM-as-judge, ground-truth recovery, execution-based correctness) and in how common criteria like “novelty” or “feasibility” are operationalized. Moreover, most existing benchmarks evaluate creativity primarily at the level of the final artifact. Since the generative process is treated as a black box, these benchmarks cannot directly assess the source or mechanism of creativity, and instead focus on observable outcomes such as novelty, plausibility, or executability. As a consequence, many evaluations rely on short-term or proxy-based metrics, leaving the dynamics of idea generation and longer-term scientific impact largely unexamined. (Ismayilzada et al., 2025b). Fourth, there is the risk of bias amplification. Both human reviewers and LLM-based judges can propagate historical inequities, undervaluing unconventional or marginalized research directions unless carefully calibrated (Gehrmann et al., 2022). Finally, a persistent scalability gap remains unresolved: while expert review provides nuance and reliability, it is expensive and prone to fatigue effects, whereas automated metrics scale efficiently but fail to capture the depth of human judgment (Thelwall, 2025).
@@
8 Ethical Concerns
While the integration of LLMs into scientific ideation substantially lowers the time and cost of hypothesis generation, experimental design, and exploratory research, it simultaneously introduces a range of ethical risks. As idea generation becomes automated and scalable, evaluation and verification must also be automated, since manual review cannot keep pace with the growing volume of outputs. This pressure is evident in contemporary peer review: NeurIPS 2025 received over 21,500 submissions, more than double its 2022 count (Foundation, 2025), far exceeding the capacity of available expert reviewers and motivating the use of automated evaluation tools such as LLM-as-a-judge (Section 7.3). However, as discussed in the previous section, reliable evaluation is fundamentally harder than generation. Evaluation must verify correctness, feasibility, evidential grounding, and must distinguish genuine creativity—combining novelty and scientific value—from fabrication, hallucination, or superficially convincing but flawed ideas. This difficulty gives rise to the risk of plausible pseudo-science, where LLM-generated ideas appear coherent and innovative yet rest on incorrect assumptions or unsound reasoning that automated evaluators struggle to detect (Li et al., 2025a). In addition, LLM-based ideation and evaluation systems may amplify existing biases in the scientific literature(Zhu et al., 2025a), preferentially favoring dominant methods, paradigms, or problem formulations while neglecting underexplored or unconventional directions, thereby reinforcing prevailing scientific dogmas. Beyond these modeling limitations, automation introduces security risks: recent work has demonstrated that LLM-based reviewers are vulnerable to prompt injection attacks (Collu et al., 2025; Lin, 2025; Zhu et al., 2025a; Theocharopoulos and others, 2025), in which adversarial text embedded in manuscripts manipulates automated reviews, highlighting how evaluation pipelines can be systematically exploited. More broadly, the use of LLMs for scientific ideation raises dual-use concerns, particularly in domains such as chemistry and biology, where generated ideas or procedures could be misused if insufficient safeguards are in place(Urbina et al., 2022). While recent studies and guidelines have begun to outline responsible practices for using AI in scientific research(Resnik and Hosseini, 2024; Knöchel et al., 2025), substantial challenges remain for governance and policymaking to adequately address verification robustness, bias, adversarial manipulation, and dual-use risks in automated scientific discovery.